# Molecular relaxation by reverse diffusion with time step prediction

### v1

@@ 3.2 Molecular relaxation performance
Figure 4: The RMSD ratios of structures relaxed with the baselines and the MoreRed variants for 20 000 non-equilibrium structures from the QM7-X test split (left) and 6500 non-equilibrium structures obtained by adding Gaussian noise over 250 forward diffusion steps to equilibrium structures from the QM7-X test split (right). The median values and the percentage of failure cases, i.e., the cases where the RMSD ratio $\geq 1$, are shown above each box plot. In Figure 5, we provide examples of relaxed structures comparing both MoreRed and the MLFF model.
@@
We assess the robustness of the methods on synthetically generated inputs by diffusing equilibrium test structures from QM7-X with $250$ forward diffusion steps. We ensure that the resulting median RMSD between the diffused configurations and the equilibrium test structures is within the range of the RMSD between the non-equilibrium structures and the equilibrium test structures from QM7-X. Figure 4a (right) shows the RMSD ratios after molecular relaxation of the diffused structures. For all MoreRed variants, the median RMSD ratio further improves compared to relaxing the QM7-X test structures, which is expected as our method is trained to denoise diffused structures. However, there is an increase in failure cases, which we attribute to more often ending up in equilibrium states different from the reference geometry. We hypothesize that this is caused by the physically less plausible deviations in the diffused structures compared to the normal mode displaced structures in QM7-X. The existence of physically less plausible deviations is supported by the results of the baseline methods, where we observe a clear deterioration in performance. MMFF94 completely fails to handle input structures perturbed with Gaussian noise and, in nearly all cases, just returns the non-equilibrium input structure, leading to a median RMSD ratio of $1$ and resulting in $99.6$% failure cases. The median RMSD ratio of GFN2-xTB as well as its percentage of failure cases are more than doubled. Most notably, the MLFF fails to get closer to the reference geometry in almost 90% of the cases. For the MLFF, this is expected as the training data distribution of non-equilibrium structures from QM7-X does not cover all the chemical space, including the Gaussian diffused inputs. Therefore, relaxation often completely fails, leading to disconnected structures even if the input structure does not appear to be overly distorted. We show an example of this in Figure 9b, with a detailed discussion in section B.3, in the supplementary information. This means that, although the MLFF uses 100 times more training data than MoreRed, it cannot easily be transferred to relax the diffused structures. In contrast, MoreRed performs well on both the diffused samples and on the non-equilibrium structures from QM7-X albeit requiring only the unlabeled equilibrium structures for training. Accordingly, the diffusion training scheme leads to a more robust method for relaxation that can be used for input structures that are obtained from different sources, e.g. different datasets, various empirical force fields, or other generative models. This also means that MoreRed will oftentimes find a reasonable structure even if the input was physically not plausible, which should be considered by practitioners using the method.
@@
Figure 5: A pairwise comparison of molecular structures from QM7-X that were relaxed with both MoreRed-AS and MLFF. In each block, from left to right, the equilibrium structure is shown (eq.), followed by the non-equilibrium structure and the structures relaxed by MoreRed-AS and MLFF, all labelled with their respective RMSD to the equilibrium structure. The first two structures follow the CPK colouring, i.e. white for hydrogen, gray for carbon, blue for nitrogen, and red for oxygen. In the subsequent two depictions, the hydrogen atoms are suppressed and the equilibrium structure (gray) and the non-equilibrium structure (white) are superimposed with MoreRed-AS and MLFF relaxed structure in green (RMSD decreased) or red (RMSD increased).
@@
For illustration, we show a series of pairwise comparisons of molecular structures from QM7-X relaxed by both MoreRed-AS and MLFF methods in Figure 5. Each panel contains a sequence of depictions, beginning with the equilibrium structure (eq.), followed by the corresponding non-equilibrium structure, and then structures relaxed using MoreRed-AS and MLFF respectively. The RMSD values, relative to the equilibrium structure, are provided for each case. For both models, we show examples that were successfully relaxed as well as failure cases. Additionally, in Figure 9 and section B.3 in the supplementary information, we provide further examples and discussion of relaxed structures using all MoreRed variants and baseline models, including MMFF94 and GFN2-xTB.
@@
Finally, we compare the energies of relaxed structures using DFT calculations. We randomly sample 100 non-equilibrium test inputs from QM7-X and compute the energy of the corresponding equilibrium reference structure as well as of the structures resulting from the relaxation of the test inputs with all methods. The energies are calculated with PBE0+MBD [102, 103] and a def2-TZVP [118] basis, using the pySCF [119, 120, 121] implementation. The results are reported in Figure 6, where we compare the deviation of the energy of the structures resulting from relaxation with the different methods from the energy of the reference equilibrium structures reported in QM7-X. Positive energy differences occur when the relaxation method yields a structure with higher energy than the reference structure, and a negative difference indicates that the molecular relaxation yields a structure with lower energy than the reference. The structures relaxed with MMFF94 mostly have significantly larger energies than the reference structures. Hence, MMFF94 is clearly outperformed by MoreRed and the other baselines, as it results in the largest structural deviations in terms of the RMSD ratio and the worst energy levels in our DFT calculations. The MLFF, on the other hand, mostly finds structures with lower energy levels compared to the reported reference structures in QM7-X. According to the previously discussed structure-label mismatch in the dataset, this can be attributed to training the MLFF on energy and force labels calculated with the PBE0+MBD, whereas the reference equilibrium structures were found with DFTB+MBD. These reported equilibrium structures still have a mean and median force magnitude greater than $5$ kcal/mol/Å when calculated with PBE0+MBD instead of DFTB+MBD, i.e. they are no minima on the PES defined by PBE0+MBD. The energy of structures relaxed with GFN2-xTB is higher than with the MLFF but still lower than the energy of the DFTB+MBD reference structures. This shows that the larger RMSD ratios found for the MLFF and GFN2-xTB stem from finding minima on a more accurate PES that is different from the one described by the QM7-X reference equilibrium structures. The energy levels of structures relaxed with MoreRed-ITP/-JT match the energy levels of the QM7-X reference structures the closest. For a majority of relaxed structures, the energy is less than 1 kcal/mol higher, i.e. within chemical accuracy. Only for MoreRed-AS, which had the smallest median RMSD ratio, we observe higher energy levels than expected in the relaxed structures. It also performs more reverse denoising steps than the other two variants most of the time, further reducing the structural deviation while underestimating the interatomic distances, which are more strongly penalized in energy calculations than in the RMSD metric.
@@
Figure 6: Deviation of the energy of structures resulting from relaxation with the different methods from the energy of the reference structures. Negative values indicate lower energy than the equilibrium reference structure. The DFT energy calculations were performed with PBE0+MBD and a def2-tzvp basis for 100 randomly sampled test structures from QM7-X.
@@
In summary, we find that MoreRed accurately captures the data manifold of equilibrium structures as it outputs relaxed structures that are close to the reference structures in both structural deviation and in terms of energy difference. Therefore, using equilibrium structures with lower energy during training may further improve MoreRed’s energy performance. Furthermore, it outperforms the classical MMFF94 method in all of our experiments and metrics. While the MLFF and GFN2-xTB find structures with larger structural deviations, the obtained structures have lower energies according to reference calculations with PBE0+MBD. This can be explained by the mismatch of computational methods employed in QM7-X to obtain equilibrium structures (DFTB+MBD) and to compute the energy and force labels (PBE0+MBD), which impedes a fair comparison. MoreRed shows improved robustness to distorted inputs compared to all baseline methods, allowing it to relax structures from different sources albeit requiring only unlabeled equilibrium structures for training. When considering all metrics together, MoreRed-JT, which uses an adaptive schedule and predicts both the time step as well as the noise for reverse diffusion with the same neural network, performs better than MoreRed-ITP/-AS and is therefore recommended.
@@
Efficiency-wise, we note that training MoreRed takes $1.45$ days on average, while the MLFF model needs more than $7$ days on NVIDIA P100. The relaxation of one single structure with MMFF94 takes $0.022$s. A single relaxation step per structure with the MLFF, GFN2-xTB, and MoreRed takes $0.02$s, $1.5$s, and $0.03$s, respectively. In contrast to the other methods that relax one structure at a time, MoreRed relaxes the structures batchwise, which could yield an even higher speed-up on tensor units. In our experiments, using batches of 128 structures results in $0.05$s per relaxation step per batch, i.e. $0.05s/128\approx 0.0004$s per relaxation step per structure. A more detailed analysis of the computation times can be found in the supplementary information D.
@@ 4 Conclusion
While not directly comparable due to the mismatch in computational methods used to create the dataset, MoreRed exhibits accurate structure relaxation performance with significantly fewer training points and reduced training time compared to machine learning force fields. The data efficiency might prove beneficial for larger systems or extensive databases where generating a sufficient amount of accurately labeled training data, especially non-equilibrium structures, is challenging. The results also revealed an inherent issue when using mismatched methods for structure relaxation and properties calculation, as observed in the QM7-X dataset. While machine learning force fields produce more accurate minima by following the PBE0+MBD-based forces, MoreRed accurately learns the data manifold of the provided DFTB+MBD-based equilibrium structures. Utilizing more accurate minima during training can further improve MoreRed’s energy performance while maintaining its data efficiency advantage.
@@ Discussion:
One might still worry that training samples drawn from two overlapping Gaussian components will deteriorate the performance of the diffusion time prediction. Indeed, such overlapping makes the prediction harder, as can be empirically seen in the right plot in Figure 7. However, with high dimensional input space, such overlapping does not significantly affect the time prediction performance for samples perturbed with small noise variance, meaning when the diffusion time step $t$ is low. Assume that there are two training molecules $\mathbf{x}_{a},\mathbf{x}_{b}$ with the Euclidean distance $r=\|\mathbf{x}_{a}-\mathbf{x}_{b}\|$, and consider the Gaussian component centered at $\mathbf{x}_{a}$, which represents its noisy versions with standard deviation $\sigma=r$. Although, in this situation, $\mathbf{x}_{b}$ lies in the high-density shell of this Gaussian component (the bump in Figure 7 left), the noisy samples of $\mathbf{x}_{a}$ are uniformly distributed all over the high $(D-1)$-dimensional shell. Therefore the probability that the Gaussian noise produces a sample close to $\mathbf{x}_{b}$ is extremely low. On the other hand, many training samples from the neighbourhood of $\mathbf{x}_{b}$ are fed to the diffusion time predictor as slightly noisy versions of $\mathbf{x}_{b}$, because of its high density. Accordingly, the network is trained to recognize the molecules close to $\mathbf{x}_{b}$ as low noise samples resulting from $\mathbf{x}_{b}$ without being disturbed by high noise samples from $\mathbf{x}_{a}$. This intuition can be mathematically confirmed by computing the density ratio between two Gaussian components around $\mathbf{x}_{b}$, i.e., $\mathcal{N}(\mathbf{x}_{b}+\bm{\varepsilon};\mathbf{x}_{b},\delta^{2}\mathbf{I})/\mathcal{N}(\mathbf{x}_{b}+\bm{\varepsilon};\mathbf{x}_{a},r^{2}\mathbf{I})$ for $\|\bm{\varepsilon}\|\sim\delta\ll r$, which is extremely high unless $D$ is very small.
To conclude, the high dimensionality of the data space enables accurate diffusion time step prediction, especially for the samples close to one of the training equilibrium molecules. This can also be observed empirically. Figure 7 (right) with orange dots shows a scatter plot of the true diffusion time vs. its prediction by our diffusion time predictor after training on the equilibrium molecules from the QM7-X dataset. We also show the performance of the diffusion time predictor trained on CIFAR10 [125] — a common image benchmark dataset — where the images have a higher dimensionality than the molecule data, as blue dots. As discussed above, the diffusion time prediction is easier when the dimension $D$ is large, and the true diffusion time, i.e., the noise level, is small.

### v2

@@ 3.2 Molecular relaxation performance
Figure 4: The RMSD ratios, i.e. the RMSD after relaxation divided by the RMSD of the non-equilibrium inputs, of structures relaxed with the baselines and the MoreRed variants, trained on the DFTB3+MBD equilibrium structures reported in QM7-X, for 20 000 non-equilibrium structures from the QM7-X test split (left) and 6500 non-equilibrium structures obtained by adding Gaussian noise over 250 forward diffusion steps to equilibrium structures from the QM7-X test split (right). The median values and the percentage of failure cases, i.e. the cases where the RMSD ratio $\geq 1$, are shown to the left of each box plot in bold. In Figure 5, we provide examples of relaxed structures comparing both MoreRed and the MLFF model.
@@
We assess the robustness of the methods on synthetically generated inputs by diffusing equilibrium test structures from QM7-X with $250$ forward diffusion steps. We ensure that the resulting median RMSD between the diffused configurations and the equilibrium test structures is within the range of the RMSD between the non-equilibrium structures and the equilibrium test structures from QM7-X. Figure 4a (right) shows the RMSD ratios after molecular relaxation of the diffused structures. For all MoreRed variants, the median RMSD ratio further improves compared to relaxing the QM7-X test structures, which is expected as our method is trained to denoise diffused structures. However, there is an increase in failure cases, which we attribute to more often ending up in equilibrium states different from the reference geometry. We hypothesize that this is caused by the physically less plausible deviations in the diffused structures compared to the normal mode displaced structures in QM7-X. The existence of physically less plausible deviations is supported by the results of the baseline methods, where we observe a clear deterioration in performance. MMFF94 completely fails to handle input structures perturbed with Gaussian noise and, in nearly all cases, just returns the non-equilibrium input structure, leading to a median RMSD ratio of $1$ and resulting in $99.6$% failure cases. The median RMSD ratio of GFN2-xTB as well as its percentage of failure cases are more than doubled. Most notably, the MLFF fails to get closer to the reference geometry in almost 90% of the cases. For the MLFF, this is expected as the training data distribution of non-equilibrium structures from QM7-X does not cover all the chemical space, including the Gaussian diffused inputs. Therefore, relaxation often completely fails, leading to disconnected structures even if the input structure does not appear to be overly distorted. We show an example of this in Figure 9b, with a detailed discussion in B.3, in the supplementary information. This means that, although the MLFF uses 100 times more training data than MoreRed, it cannot easily be transferred to relax the diffused structures. In contrast, MoreRed performs well on both the diffused samples and on the non-equilibrium structures from QM7-X albeit requiring only the unlabeled equilibrium structures for training. Accordingly, the diffusion training scheme leads to a more robust method for relaxation that can be used for input structures that are obtained from different sources, e.g. different datasets, various empirical FFs, or other generative models. This also means that MoreRed will oftentimes find a reasonable structure even if the input was physically not plausible, which should be considered by practitioners using the method.
@@
Figure 5: A pairwise comparison of molecular structures from QM7-X that were relaxed with both MoreRed-AS (MR-AS) and MLFF. In each block, from left to right, the equilibrium structure is shown (eq.), followed by the non-equilibrium structure (non-eq.) and the structures relaxed by MoreRed-AS and MLFF, all labelled with their respective RMSD to the equilibrium structure in angstrom (Å). The first two structures follow the CPK colouring, i.e. white for hydrogen, gray for carbon, blue for nitrogen, and red for oxygen. In the subsequent two depictions, the equilibrium structure (CPK) is superimposed with MoreRed-AS and MLFF relaxed structure. Structures with an decreased RMSD value relative to the non-equilibrium structure are shown in green, while failure cases with a higher RMSD compared to the initial structure are highlighted in salmon
@@
For illustration, we show a series of pairwise comparisons of molecular structures from QM7-X relaxed by both MoreRed-AS and MLFF methods in Figure 5. Each panel contains a sequence of depictions, beginning with the equilibrium structure (eq.), followed by the corresponding non-equilibrium structure, and then structures relaxed using MoreRed-AS and MLFF respectively. The RMSD values, relative to the equilibrium structure, are provided for each case. For both models, we show examples that were successfully relaxed as well as failure cases. Additionally, in Figure 9 and B.3 in the supplementary information, we provide further examples and discussion of relaxed structures using all MoreRed variants and baseline models, including MMFF94 and GFN2-xTB.
@@
Finally, we compare the energies of relaxed structures using DFT calculations. We randomly sample 100 non-equilibrium test inputs from QM7-X and compute the energy of the corresponding equilibrium reference structure as well as of the structures resulting from the relaxation of the test inputs with all methods. The energies are calculated with PBE0+MBD [109, 110] and a def2-TZVP [125] basis, using the PySCF [126, 127, 128] implementation. The results are reported in Figure 6, where we compare the deviation of the energy of the structures resulting from relaxation with the different methods from the energy of the reference equilibrium structures reported in QM7-X. Positive energy differences occur when the relaxation method yields a structure with higher energy than the reference structure, and a negative difference indicates that the molecular relaxation yields a structure with lower energy than the reference. The structures relaxed with MMFF94 mostly have significantly larger energies than the reference structures. Hence, MMFF94 is clearly outperformed by MoreRed and the other baselines, as it results in the largest structural deviations in terms of the RMSD ratio and the worst energy levels in our DFT calculations. The MLFF, on the other hand, mostly finds structures with lower energy levels compared to the reported reference structures in QM7-X. According to the previously discussed structure-label mismatch in the dataset, this can be attributed to training the MLFF on energy and force labels calculated with the more accurate PBE0+MBD, whereas the reference equilibrium structures were found with DFTB+MBD. These reported equilibrium structures still have a mean and median force magnitude greater than $5$ kcal/mol/Å when calculated with PBE0+MBD instead of DFTB+MBD, i.e. they are no minima on the PES defined by PBE0+MBD. The energy of structures relaxed with GFN2-xTB is higher than with the MLFF but still lower than the energy of the DFTB+MBD reference structures. This shows that the larger RMSD ratios found for the MLFF and GFN2-xTB stem from finding minima on a more accurate PES that is different from the one described by the QM7-X reference equilibrium structures. The energy levels, indicated by the darker colors in Figure 6, of structures relaxed with MoreRed-ITP/-JT, trained on the reference DFTB3+MBD minima from QM7-X, match the energy levels of these reference structures the closest. For a majority of relaxed structures, the energy is less than 1 kcal/mol higher, i.e. within chemical accuracy. Only for MoreRed-AS, which had the smallest median RMSD ratio, we observe higher energy levels than expected. It also performs more reverse denoising steps than the other two variants most of the time, further reducing the structural deviation while underestimating the interatomic distances, which are more strongly penalized in energy calculations than in the RMSD metric. Overall, we observe a mismatch between the RMSD results and the DFT energies of the MLFF and all MoreRed variants that were trained on equilibrium structures from QM7-X (darker colors in Figure 6), where methods having lower RMSD ratios result in higher energies and vice versa. Therefore, to further investigate this mismatch and ensure a fair comparison, we further relaxed the DFTB+MBD equilibrium structures of QM7-X using the MLFF to minimize the forces and consequently the energies. We then retrained all MoreRed variants on this MLFF-relaxed dataset and used them to relax the non-equilibrium test structures from QM7-X. The resulting DFT energies, represented with light-colored boxes in Figure 6, show that MoreRed models, trained on the MLFF-relaxed structures, achieve much lower PBE0+MBD energies compared to those trained on the original DFTB+MBD minima. However, they exhibit higher RMSD ratios, as demonstrated in Figure 10 in the supplementary information. This outcome aligns with the results achieved by the MLFF and supports the notion that the RMSD-energy mismatch is related to the structure-label mismatch. Specifically, the MLFF finds lower PBE0+MBD energy minima that differ from the reference DFTB+MBD minima reported in QM7-X, resulting in higher RMSD ratios. Further details on this experiment are provided in B.4 of the supplementary information.
@@
Figure 6: Deviation of the energy of structures resulting from relaxation with the different methods from the energy of the reference structures from QM7-X. Negative values indicate lower energy than the equilibrium reference structure. The DFT energy calculations were performed with PBE0+MBD and a def2-TZVP basis for 100 randomly sampled test structures from QM7-X. Each MoreRed variant has two versions: the darker color represents the model trained on the original DFTB3+MBD equilibrium structures from QM7-X, which have non-negligible PBE0+MBD residual forces, and uses the same relaxation results used for Figure 4. On the other hand, the lighter color represents the model trained on equilibrium structures resulting from further relaxation of the QM7-X reference structures to minimize the residual forces, using the MLFF trained on the more accurate PBE0+MBD force labels.
@@
In summary, we find that MoreRed accurately captures the data manifold of equilibrium structures as it outputs relaxed structures that are close to the reference structures in both structural deviation and in terms of energy difference. Moreover, using equilibrium structures with lower energy during training further improves MoreRed’s energy performance. MoreRed outperforms the classical MMFF94 method in all of our experiments and metrics. While the MLFF and GFN2-xTB find structures with larger structural deviations, the obtained structures have lower energies according to reference calculations with PBE0+MBD. This was explained by the mismatch of computational methods employed in QM7-X to obtain equilibrium structures (DFTB+MBD) and to compute the energy and force labels (PBE0+MBD). Besides, MoreRed shows improved robustness to distorted inputs compared to all baseline methods, allowing it to relax structures from different sources albeit requiring only unlabeled equilibrium structures for training. When considering all metrics together, MoreRed-JT, which uses an adaptive schedule and predicts both the time step as well as the noise for reverse diffusion with the same neural network, performs better than MoreRed-ITP/-AS and is therefore recommended.
@@
Efficiency-wise, we note that training MoreRed takes $1.45$ days on average, while the MLFF model needs more than $7$ days on NVIDIA P100. The relaxation of one single structure with MMFF94 takes $0.022$s. A single relaxation step per structure with the MLFF, GFN2-xTB, and MoreRed takes $0.02$s, $1.5$s, and $0.03$s, respectively. In contrast to the other methods that relax one structure at a time, MoreRed relaxes the structures batchwise, which could yield an even higher speed-up on tensor units. In our experiments, using batches of 128 structures results in $0.05$s per relaxation step per batch, i.e. $0.05s/128\approx 0.0004$s per relaxation step per structure. A more detailed analysis of the computation times can be found in the supplementary information D. Moreover, the effect of the size of the molecule on the performance and efficiency of MoreRed is studied in E.
@@ 4 Conclusion
While not directly comparable due to the mismatch in computational methods used to create the dataset, MoreRed exhibits accurate structure relaxation performance with significantly fewer training points and reduced training time compared to machine learning FFs. The data efficiency might prove beneficial for larger systems or extensive databases where generating a sufficient amount of accurately labeled training data, especially non-equilibrium structures, is challenging. The results also revealed an inherent issue when using mismatched methods for structure relaxation and properties calculation, as observed in the QM7-X dataset. While machine learning FFs produce more accurate minima by following the PBE0+MBD-based forces, MoreRed accurately learns the data manifold of the provided DFTB+MBD-based equilibrium structures. Utilizing more accurate minima during training can further improve MoreRed’s energy performance while maintaining its data efficiency advantage.
@@ Discussion:
One might still worry that training samples drawn from two overlapping Gaussian components will deteriorate the performance of the diffusion time prediction. Indeed, such overlapping makes the prediction harder, as can be empirically seen in the right plot in Figure 7. However, with high dimensional input space, such overlapping does not significantly affect the time prediction performance for samples perturbed with small noise variance, meaning when the diffusion time step $t$ is low. Assume that there are two training molecules $\mathbf{x}_{a},\mathbf{x}_{b}$ with the Euclidean distance $r=\|\mathbf{x}_{a}-\mathbf{x}_{b}\|$, and consider the Gaussian component centered at $\mathbf{x}_{a}$, which represents its noisy versions with standard deviation $\sigma=r$. Although, in this situation, $\mathbf{x}_{b}$ lies in the high-density shell of this Gaussian component (the bump in Figure 7 left), the noisy samples of $\mathbf{x}_{a}$ are uniformly distributed all over the high $(D-1)$-dimensional shell. Therefore the probability that the Gaussian noise produces a sample close to $\mathbf{x}_{b}$ is extremely low. On the other hand, many training samples from the neighbourhood of $\mathbf{x}_{b}$ are fed to the diffusion time predictor as slightly noisy versions of $\mathbf{x}_{b}$, because of its high density. Accordingly, the network is trained to recognize the molecules close to $\mathbf{x}_{b}$ as low noise samples resulting from $\mathbf{x}_{b}$ without being disturbed by high noise samples from $\mathbf{x}_{a}$. This intuition can be mathematically confirmed by computing the density ratio between two Gaussian components around $\mathbf{x}_{b}$, i.e. $\mathcal{N}(\mathbf{x}_{b}+\bm{\varepsilon};\mathbf{x}_{b},\delta^{2}\mathbf{I})/\mathcal{N}(\mathbf{x}_{b}+\bm{\varepsilon};\mathbf{x}_{a},r^{2}\mathbf{I})$ for $\|\bm{\varepsilon}\|\sim\delta\ll r$, which is extremely high unless $D$ is very small.
To conclude, the high dimensionality of the data space enables accurate diffusion time step prediction, especially for the samples close to one of the training equilibrium molecules. This can also be observed empirically. Figure 7 (right) with orange dots shows a scatter plot of the true diffusion time vs. its prediction by our diffusion time predictor after training on the equilibrium molecules from the QM7-X dataset. We also show the performance of the diffusion time predictor trained on CIFAR10 [132] — a common image benchmark dataset — where the images have a higher dimensionality than the molecule data, as blue dots. As discussed above, the diffusion time prediction is easier when the dimension $D$ is large, and the true diffusion time, i.e. the noise level, is small.